Help me solve this equation. The equation has to be solved (a)symbolically (b)graphically, and (c) numerically
4x+y=6
x-4=14
Please help, this has 23 diffrent steps and I am at a loss!

Answer by Alan3354(31538)

4x+y=6
x-4=14 --> x = 18
----------
4*18 + y = 6
y = -66
Not so many.